Presented here are summarized data from the balance sheets and income statements of Wiper Inc.:

WIPER INC.
Condensed Balance Sheets
December 31, 2020, 2019, 2018
(in millions)
2020 2019 2018
Current assets $ 793 $ 1,021 $ 883
Other assets 2,428 1,935 1,734
Total assets $ 3,221 $ 2,956 $ 2,617
Current liabilities $ 592 $ 845 $ 742
Long-term liabilities 1,607 1,070 942
Stockholders’ equity 1,022 1,041 933
Total liabilities and stockholders' equity $ 3,221 $ 2,956 $ 2,617
WIPER INC.
Selected Income Statement and Other Data
For the year Ended December 31, 2020 and 2019
(in millions)
2020 2019
Income statement data:
Sales $ 3,065 $ 2,928
Operating income 311 325
Interest expense 99 80
Net income 236 231
Other data:
Average number of common shares outstanding 42.8 48.2
Total dividends paid $ 65.0 $ 53.8


Required:

  1. Calculate return on investment, based on net income and average total assets, for 2020 and 2019.
  2. Calculate return on equity for 2020 and 2019.
  3. Calculate working capital and the current ratio for each of the past three years.
  4. Calculate earnings per share for 2020 and 2019.
  5. If Wiper's stock had a price/earnings ratio of 12 at the end of 2020, what was the market price of the stock?
  6. Calculate the cash dividend per share for 2020 and the dividend yield based on the market price calculated in part e.
  7. Calculate the dividend payout ratio for 2020.
  8. Assume that accounts receivable at December 31, 2020, totaled $324 million. Calculate the number of days' sales in receivables at that date.
  9. Calculate Wiper's debt ratio and debt/equity ratio at December 31, 2020 and 2019.
  10. Calculate the times interest earned ratio for 2020 and 2019.
Answer #1

Values are in millions,

a)  Return on investment for 2020, 2019

Return on Investment = Net income / Average of total assets

Particulars 2020 2019
1. Net Income 236 231
A) Assets in the beginning 2956 2617
B) Assets in the end 3221 2956
2. Average of total assets[(A+B)/2] 3089 2787
Return on Investment (1/2) 7.64% 8.29%

b)  Return on equity for 2020, 2019

Return on equity = Net income/ Average Stockholders' equity

Particulars 2020 2019
1. Net income 236 231
a) Equity in beginning 1041 933
b) Equity at the end 1022 1041
2. Average of Shareholders equity[(a+b)/2] 1032 987
Return on equity(1/2) 22.87% 23.4%

c) Calculation of working capital and current ratio for 2018, 2019, 2020

Working capital = Total Current assets - Total Current Liabilities

Current ratio = Current assets/Current liabilities

Particulars 2020 2019 2018
1.Total Current assets 793 1021 883
2.Total Current liabilities 592 845 742
Working capital (1-2) 201 176 141
Current ratio (1/2) 1.34 1.21 1.19

d) Calculation of Earnings per share for 2020 and 2019

Earnings per share is the earnings made per share

Earnings per share = Net income / Average number of Shares outstanding

Particulars 2020 2019
1. Net Income 236 231
2. Average number of Shares outstanding 42.8 48.2
Earnings per share (1/2) $5.51 $4.79

e) Market price of the stock for the year 2020

Price earnings ratio = 12%

Earnings per share for 2020 = $5.51

Market price of stock = price earning ratio * Earnings per share

= 12 * 5.51 = $ 66.12

f) Cash dividend per share for the year 2020 and dividend yield

Total dividends paid = $65

Shares outstanding = 42.8

Cash dividend per share = Total dividends paid/average shares outstanding

= 65/42.8 = $ 1.52 per share

Dividend yield based on market price = (Dividend per share/Market price of the stock)*100

= (1.52/ 66.12)*100

= 2.30%

g) Dividend payout ratio for 2020

Dividend payout ratio = Dividend paid per share/Net income

= 65/236 = 0.28 times

h) Calculation of number of days'sales in receivables at that date

Average sales = $3065/365 = $8.40

Accounts receivables = $324

Number of days sales in receivables = Average Accounts receivables / Average sales

= 324/8.4 = 38.57 days

i) Wiper's debt ratio and debt/equity ratio at Dec 2020, 2019

Debt ratio = Total liabilities / total assets

Debt equity ratio = Total liabilities / Total equity

Particulars 2020 2019
1. Current liabilities 592 845
2.Long term liabilities 1607 1070
a) Total liabilities (1+2) 2199 1915
b) Total assets 3221 2956
c) Total equity 1022 1041
Debt ratio (a/b) 68.27% 64.78%
Debt equity ratio (a/c) 2.15 1.84

j) Time interest earned ratio for 2020 and 2019

Time interest earned ratio = Earnings before interest and taxes/interest expense

Particulars 2020 2019
1. Net income 236 231
2. Interest 99 80
3. Tax paid 0 0
4. Eanings before interest and tax (1+2+3) 335 311
Time interest earned (4/2) 3.38 3.89